# Diffusivity-Free Turbulence in Tabletop Rotating Rayleigh-Bénard Convection Experiments

### v1

@@
Diffusivity-Free Turbulence in Tabletop Rotating Rayleigh-Bénard Convection Experiments
Jewel A. Abbate (jewelabbate@ucla.edu) Yufan Xu Tobias Vogt Susanne Horn Keith Julien Jonathan M. Aurnou
journal:
[…]
@@
Convection in planets and stars is predicted to occur in the “ultimate regime” of diffusivity-free, rapidly rotating turbulence, in which flows are characteristically unaffected by viscous and thermal diffusion. Boundary layer diffusion, however, has historically hindered experimental study of this regime. Here, we utilize the boundary-independent oscillatory thermal-inertial mode of rotating convection to realize the diffusivity-free scaling in liquid metal laboratory experiments. This oscillatory style of convection arises in rotating liquid metals (low Prandtl number fluids) and is driven by the temperature gradient in the fluid bulk, thus remaining independent of diffusive boundary dynamics. We triply verify the existence of the diffusivity-free regime via measurements of heat transfer efficiency $\boldsymbol{Nu}$, dimensionless flow velocities $\boldsymbol{Re}$, and internal temperature anomalies $\boldsymbol{\theta}$, all of which are in quantitative agreement with planar asymptotically-reduced models. Achieving the theoretical diffusivity-free scalings in desktop-sized laboratory experiments provides the validation necessary to extrapolate and predict the convective flows in remote geophysical and astrophysical systems.
@@
Diffusivity-free (DF) turbulent convection is theorized to occur in the convection zones of stars and the fluid cores of planets, where it drives vortical flows, jets, and dynamo action that sustains global-scale magnetic fields (Moffatt and Dormy, 2019). The DF regime is characterized by strong turbulence dominated by fluid inertia, largely unaffected by the fluid’s micro-scale viscosity ($\nu$) and thermal diffusivity ($\kappa$). For decades, researchers in the fluid physics and geophysics communities have sought to validate this theoretical regime through measurements of heat transfer efficiency in numerical simulations and laboratory experiments (e.g., Ahlers et al., 2009; Ecke and Shishkina, 2023). It remains an essential question whether the theoretical DF regime can be attained in a laboratory setting and be extrapolated to natural settings.
@@
The canonical system for studying rotationally-constrained turbulent flows is the rotating Rayleigh-Bénard convection (RRBC) model (Ecke and Shishkina, 2023). In RRBC, a fluid layer confined between two parallel no-slip boundaries is heated from below and cooled from above while rotating about its central vertical axis. Scaling laws that relate global heat transport (Nusselt number, $Nu$) and bulk flow velocity (Reynolds number, $Re$) to the buoyancy forcing (Rayleigh number, $Ra$), global rotation period (Ekman number, $E$), and material properties (Prandtl number, $Pr$) enable the extrapolation of laboratory-numerical findings to more extreme geophysical and astrophysical scales (Kraichnan, 1962; Spiegel, 1971; Stevenson, 1979; Christensen and Aubert, 2006; Plumley and Julien, 2019; Abbate and Aurnou, 2023). (See Table 1 for parameter definitions.)
@@
[…]
@@
Figure 1: Rotating laboratory setup. (a) Image of the RoMag device at UCLA with the $\Gamma\simeq 1/2$ tank in place. Heat is supplied to the copper base (‘heating block’) by a heating pad and extracted from the copper lid (‘cooling block’) by a heat exchanger routed to a recirculating chiller. The cell is filled with liquid gallium and wrapped in layers of insulation (not pictured). Red and black visible wires are connected to thermistors and Doppler transducer probes. (b) Schematic of measurement locations. White lines at the cell boundaries represent thermistors embedded within the heating and cooling blocks used to measure $\Delta T$. Black dots indicate the location of five internal thermistors used to measure temperature fluctuations, $\theta$. Orange lines indicate the 1D path sensed by Doppler transducers used to measure vertical flow velocity, $u_{z}$. (c) Example time series data for a turbulent case in the oscillatory regime. The top panel shows thermistor data, where shades of pink (blue) indicate bottom (top) boundary temperatures and shades of gray indicate internal temperatures. The bottom panel shows velocimetry data, where pink (blue) regions indicate upward (downward) flow. The middle 50% of the chord is used to calculate the rms-velocity, which is demarcated with dotted lines here and in panel (b). The case shown corresponds to parameters: $Ra\simeq 10^{9}$, $E\simeq 8\times 10^{-7}$, $Ro_{c}\simeq 0.16$, $Ro_{z}\simeq 0.01$, and $Re_{z}\simeq 2\times 10^{4}$.
The scaling predictions for the DF regime are derived by enforcing the constraint that thermal and viscous diffusion have no effect on the total heat transported through the system nor the internal flow velocities, and assuming that the system’s angular rotation rate goes to infinity ($\Omega\rightarrow\infty$) (Julien et al., 2012). The global heat transport prediction for rotating DF convection follows
@@
where $C_{J}$ is a constant prefactor. This value was determined by Julien et al. (2012) to be $C_{J}\approx 1/25$ in doubly-periodic planar models of the asymptotically reduced RRBC equations. Scaling predictions for the internal quantities are determined from a torque balance between the inertial and buoyancy terms of the vorticity equation describing rotating flows, neglecting viscous diffusion (e.g., Ingersoll and Pollard, 1982). Assuming any scaling coefficients other than $C_{J}$ are order unity (see Methods), the prediction for the internal flow velocity is
@@
where $\theta$ is the root-mean-squared temperature perturbation, and $Ro_{c}=(RaE^{2}Pr^{-1})^{1/2}$ is the convective Rossby number (Abbate and Aurnou, 2023, and see Methods). These scalings have yet to be unambiguously found in any closed cell RRBC laboratory setup, thus raising the question of their applicability to any natural enclosed system (cf. Soderlund et al., 2014).
@@
The fundamental barrier to achieving diffusion-free convection in RRBC lies in the physical boundaries through which heat is supplied and extracted. Boundary layers throttle energy transfer through the system. In most RRBC cases, the primary mode of convective instability is that of steady convection, which is thermally controlled by the conditions at the boundaries (King and Aurnou, 2012; Julien et al., 2016). The critical value of the Rayleigh number for which steady convective motions first develop (denoted as $Ra_{S}$, see Methods) is determined from a balance between friction and the Coriolis and pressure gradient forces, such that the steady mode is considered a viscous convective instability (Zhang and Liao, 2017; Aurnou et al., 2018). Consequently, heat transfer in typical steady RRBC is dominated by boundary layer diffusion (e.g., Abbate and Aurnou, 2023; Oliver et al., 2023), which impedes the system from reaching the asymptotic DF scaling limit.
@@
Lateral boundaries in RRBC give rise to additional sidewall-attached convective instabilities. These “wall modes” manifest as waves that travel azimuthally around the periphery of the tank, and are also a viscous instability (Ecke et al., 1992). Between contributions from steady and wall mode convection, RRBC heat transfer is often strongly influenced by diffusive effects.
@@
In order to mitigate boundary contributions in convection studies, researchers have employed innovative techniques to bypass the boundaries altogether. Numerically, this has been achieved through the use of ghost layers or forced decoupling of boundary layers (Barker et al., 2014; Zou and Yang, 2021). Experimentally, radiative heat deposition directly into the fluid bulk (Lepot et al., 2018; Bouillaut et al., 2021) has recently found the asymptotic heat transfer scaling exponent, but a very different value of the prefactor than that of Julien et al. (2012) due to the differing means of thermal energy injection. Their results may apply to the rotating convective turbulence present in gas planets and stars; however, it is uncertain that they apply to terrestrial fluid environments where convection is still driven by fluxing heat through finite, solid boundaries. Thus, bounded RRBC systems have yet to yield unambiguous asymptotic scalings, raising the dual concerns $i$) that standard laboratory-scale devices may be unable to realistically simulate geophysical turbulence or $ii$) that DF scalings do not exist in natural convection enclosed by solid boundaries (Favier and Knobloch, 2020; de Wit et al., 2020; Zhang et al., 2021; Lu et al., 2021; Madonia et al., 2023).
@@
To address these questions, we utilize the thermal-inertial oscillatory mode of convection that exists in low $Pr$ liquid metal RRBC (Chandrasekhar, 1961; Zhang and Roberts, 1997; Zhang and Liao, 2009; Horn and Schmid, 2017; Aurnou et al., 2018). Contrary to steady convection, which is subject to thermal boundary conditions, oscillatory convection is driven entirely by the internal temperature field. Zhang and Roberts (1997) showed that oscillatory convection is described by a thermal-inertial wave, arising from a balance between inertia and Coriolis acceleration, with no dependence on the nature of the boundaries or diffusive effects at dominant order. This thermally-driven inertial instability arises exclusively in rotating low $Pr<0.68$ fluids, like liquid metals and plasmas. The critical $Ra$ for which oscillatory modes onset (denoted as $Ra_{O}$, see Methods) occurs at lower $Ra$ than that for steady convection ($Ra_{S}$), such that thermal-inertial oscillations are considered the preferred mode of convection when the Prandtl number is sufficiently small (Zhang and Liao, 2017). Here, by isolating the oscillatory mode in rotating liquid gallium ($Pr\approx 0.027$) convection experiments, we show that the inertial asymptotic DF scalings (1) - (3) can be realized in a desktop-sized RRBC cell, and are in good quantitative agreement with Julien et al. (2012)’s asymptotically-reduced RRBC models, further validating both laboratory and multi-scale asymptotic methods.
@@
Figure 2: Rotating convection survey measurements. (a) Global heat transport efficiency, $Nu$, and (b) vertical non-dimensional bulk velocity, $Re_{z}$, as a function of thermal buoyancy forcing, $Ra$. Symbol color denotes the Ekman number, $E$, where red (lowest $E$) indicates the strongest Coriolis forcing and purple (highest $E$) indicates weaker Coriolis forcing. Symbol shape denotes the cell aspect ratio, $\Gamma=D/H$. Laboratory results from this study ($\Gamma\simeq 1/2$) are marked as circles, those from King and Aurnou (2013) ($\Gamma\simeq 1$) are marked as triangles, and those from Vogt et al. (2021) ($\Gamma\simeq 2$) are marked as squares. Hollow points indicate cases in which the steady mode of convection has onset ($Ra\geq Ra_{S}$), which is used here as a transition point between inertial and steady convection regimes. The solid black line in each panel shows the non-rotating empirically-determined scaling for the $\Gamma\simeq 1/2$ cell from this study for reference. A plot of the non-rotating RBC data used to generate these fits is shown in Supplementary Fig. 5. The full data set can be found in Supplementary Table LABEL:tab:data.
@@
UCLA’s RoMag device is used to conduct our experiments. The setup (Fig. 1a and Supplementary Fig. 4a) features a right cylindrical convection cell with a height of $H=40.0$ cm and a diameter of $D=19.9$ cm, yielding an aspect ratio of $\Gamma=D/H\simeq 1/2$. The cell is constructed with a stainless steel sidewall positioned between two copper end blocks. Heat is supplied through a basal heating pad, while a heat exchanger plate, connected to a re-circulating chiller, removes heat at the top. After equilibration, both the top and bottom copper end blocks are maintained at fixed temperatures for the duration of the experiment. The convection cell is affixed to a rotating platform that spins the cell about its vertical axis. We carry out a systematic survey, varying the Rayleigh and Ekman numbers by approximately two orders of magnitude each ($8\times 10^{6}<Ra<1\times 10^{9}$ and $3\times 10^{-7}<E<1\times 10^{-5}$).
@@
For each experiment, we acquire measurements of temperature, heat flux, and velocity. Thermistors embedded within the top and bottom thermal blocks measure the vertical temperature gradient, $\Delta T$ (see white lines in Fig. 1b and pink/blue time series lines in Fig. 1c). Five internal thermistors are positioned within the bulk fluid to measure the internal temperature fluctuations (see black lines in Fig. 1b and black/gray time series lines in Fig. 1c). The standard deviation of each internal thermistor time series is calculated to determine the local perturbation at each location, which are then RMS-averaged to yield $\theta$. Vertical flow velocity, $u_{z}$, is measured using an ultrasonic Doppler velocimeter (UDV). A Doppler transducer probe is positioned in contact with the interior fluid, oriented along the vertical axis of the cylinder. Natural impurities in the gallium act as particles that the UDV can acoustically sense (Wang and Kelley, 2021). Velocities are recorded at approximately 2 mm increments across the entire vertical length of the cell at a sampling rate of approximately 10 Hz. The characteristic $u_{z}$ velocity is then calculated as a time- and space-RMS, using the middle 50% of the cell height (see orange lines and light orange area shaded in Fig. 1b and the Hovmöller diagram in 1c). The equations used for averaging are included in Methods.
@@
We compare our $\Gamma\simeq 1/2$ convection cell results to those of two other aspect ratios from prior studies performed using the same experimental device. King and Aurnou (2013) conducted rotating convection experiments in liquid metal using a $\Gamma\simeq 1$ cell, measuring global heat transport efficiency, $Nu$, with their data reported in King and Aurnou (2015). Vogt et al. (2021) conducted experiments using a $\Gamma\simeq 2$ cell, measuring the Nusselt number, $Nu$, vertical Reynolds number, $Re_{z}$, and internal temperature fluctuations, $\theta$. We additionally compare our findings to the asymptotically-reduced planar models of Julien et al. (2012) by incorporating their best-fit prefactor $C_{J}=1/25$ in the heat transfer prediction (1). We elect to carry this prefactor through the asymptotic predictions for the Reynolds number (2) and thermal anomalies (3), as shown in Methods.
@@
Figure 3: Laboratory scaling results. (a) The convective heat flux relative to conductive heat flux, $Nu-1$, (b) the normalized internal temperature perturbation, $\theta/\Delta T$, and (c) the measured vertical Reynolds number, $Re_{z}$, each shown as a function of the asymptotically reduced scaling result from Julien et al. (2012) (see Methods for details). Solid black lines indicate the asymptotically reduced predictions. Dotted lines show the best-fit power law results to each $\Gamma$ data set, where the hollow points (indicating $Ra/Ra_{S}\geq 1$) are excluded from the fit. The inset plots display the best-fit prefactors as hollow symbols and the best fit scaling exponents as solid symbols, with solid horizontal lines marking the asymptotic predictions. The best-fit scaling relations are each listed in Table 2. Across all three measured quantities, the scaling approaches the asymptotic result as the aspect ratio, $\Gamma$, increases. Other markers are the same as those in Fig. 2.
@@
Fig. 2 presents the measured heat transfer ($Nu$) and vertical velocities ($Re_{z}$) across all three aspect ratios considered, plotted as a function of the buoyancy forcing ($Ra$). Rotating data are indicated by colored markers, while a solid black line represents the empirically determined non-rotating scaling relation (data shown in Supplementary Fig. 5). Solid colored points indicate data within the oscillatory convection regime. Hollow points represent experiments where steady convection has additionally onset, following $Ra\geq Ra_{S}=(8.70-9.63E^{1/6})E^{-4/3}$ (Niiler and Bisshopp, 1965; Kunnen, 2021). These cases are not expected to follow the inertial DF scaling predicted for the oscillatory mode (see Supplementary Fig. 6) and are therefore not included in the scaling analysis. Additionally, in low Prandtl number fluids, a weakly nonlinear regime is present near onset ($1\leq Ra/Ra_{O}\lesssim 1.5$) (Jones et al., 1976; Chiffaudel et al., 1987). Thus, only measurements where $Ra/Ra_{O}>1.5$ are considered here. Notably, all cases exhibit high Reynolds numbers with $300\lesssim Re\lesssim 3000$. Even near convective onset ($Nu\leq 2$), the Reynolds numbers reach as high as $Re\simeq 10^{3}$.
Fig. 3 compares our measurements with the theoretical DF predictions formulated with Julien et al. (2012)’s $C_{J}=1/25$ prefactor. The data is presented as colored markers, with a solid black line indicating the asymptotically reduced predictions. Power law fits to the oscillatory data (solid points) are shown as dotted lines of the same color. Inset plots display the best-fit values, and Table 2 lists the specific best-fit power law relations for each dataset. Pre-factors and exponents of unity, which are marked by a solid black horizontal line in the inset plots, correspond to the asymptotically reduced result. The oscillatory data is in good quantitative agreement with this prediction. Hollow points (indicating $Ra>Ra_{S}$) clearly deviate from the diffusion-free scaling. Thus, the thermal-inertial modes follow the non-diffusive scalings, whilst the inclusion of bulk stationary modes truncates this behavior.
The heat transfer efficiency ($Nu$) results in Fig. 3a show a close fit between the asymptotically reduced and $\Gamma\simeq 2$ cell data. Decreasing the aspect ratio likely introduces strong wall mode effects that enhance global heat transfer efficiency (Favier and Knobloch, 2020; de Wit et al., 2020; Zhang et al., 2021; Lu et al., 2021), causing the slope to deviate from the asymptotic scaling. The inset plot, which shows the best-fit values, reveals a clear trend toward the asymptotic limit as the aspect ratio increases for both the scaling exponent and the prefactor. Thus, the wider aspect ratio cell more closely replicates infinite plane layer behavior than the narrower aspect ratio cells, making it a better fit to the asymptotically reduced model result.
The thermal anomaly ($\theta$) results shown in Fig. 3b exhibit excellent agreement with theory across both aspect ratios ($\Gamma\simeq 1/2$, $\Gamma\simeq 2$). The inset plots show scaling exponents near unity for both datasets. Unlike global heat transfer, internal fluctuations appear to be insensitive to wall flow behaviors such that they exclusively follow the thermal-inertial bulk scaling trends regardless of aspect ratio.
The vertical velocity $Re_{z}$ results shown in Fig. 3c also strongly align with diffusion-free theory, as they are similarly unaffected by wall modes. Overall, we find simultaneous agreement with diffusion-free predictions across all three independently measured quantities in this study: thermal anomalies in the bulk, velocities in the bulk, and global heat transfer. Further, we find increasing fidelity of this agreement in higher aspect ratio containers.
These experiments show that diffusion-free physics can exist even in relatively small, closed containers of rotating fluid. Of specific note, we find that rotating low $Pr$ liquid metal convection in a short, wide aspect ratio tank ($\Gamma\simeq 2$) robustly agrees with the varied $Pr$ results of the asymptotically-reduced planar modeling study of Julien et al. (2012). This work thus verifies the $C_{J}$ constant found by Julien et al. (2012) to hold for DF planar systems with no-slip top and bottom boundaries. In our laboratory setup, we utilize the oscillatory regime of low $Pr$ rotating convection to achieve this goal. However, Julien et al. (2012) find that the scaling can hold for a range of $Pr$ assuming the criteria of rotationally-constrained turbulence are met ($Ro\ll 1\ll Re$), as expected in planetary interior flows (Holme et al., 2015; Soderlund et al., 2024) and in many stellar settings as well (Vasil et al., 2021).
The laboratory-theoretical agreement crucially validates application of DF theory to natural settings. Here we consider Earth’s outer core as an example. Scalings (1) - (3) are cast in terms of the superadiabatic temperature difference, $\Delta T$, where a known Rayleigh number or convective Rossby number is used to predict an unknown heat transfer, velocity, and thermal anomaly. However, the Rayleigh number is rarely known in geophysical or astrophysical applications; instead, a heat flux or flow velocity is typically estimated from observational data. Specifically for Earth’s outer core, flow velocities are estimated from geomagnetic secular variation to be $u\simeq 5\times 10^{-4}$ m/s (Holme et al., 2015), which gives a Reynolds number of $Re\simeq 10^{9}$. The diffusion-free Reynolds number equation (1) can then be converted to a flux-based format and rearranged to solve for the Rayleigh number as an unknown quantity:
$Ra=C_{J}^{-2/5}ReE^{-1}Pr.$ (4)
Using (4) with $C_{J}=1/25$, $Re\simeq 10^{9}$, $E\simeq 10^{-15}$, and $Pr\simeq 0.1$ for Earth’s outer core, the Rayleigh number is calculated to be $Ra\simeq 4\times 10^{23}$. The Rayleigh number estimate can then be used to determine the superadiabatic temperature differential $\Delta T$ in the core following $\Delta T=Ra\cdot\nu\kappa/(\alpha g_{cmb}H^{3})\simeq 3$ mK, where $\nu\simeq 10^{-6}$ m2/s, $\kappa\simeq 10^{-5}$ m2/s, $\alpha\simeq 1.2\times 10^{-5}$ 1/K, $g_{cmb}\simeq 10.68$ m/s, and $H\simeq 2255$ km (Olson, 2015). Using (3), the temperature anomaly is calculated as $\theta=(1/25)^{3/5}(RaE^{2}Pr^{-1})^{1/2}\Delta T\simeq 10^{-6}$ K.
Equation (4) can additionally be coupled with the heat transfer prediction (1) to estimate the power driving thermal convection in the core (neglecting compositional effects). We multiply both sides of (1) by the Rayleigh number to solve for the flux Rayleigh number, $Ra_{F}=Ra(Nu-1)=\alpha gq_{conv}H^{4}/(\nu\kappa k)$, where $q_{conv}$ is the convective heat flux, leading to the expression:
$Ra_{F}=C_{J}Ra^{5/2}E^{2}Pr^{-1/2}.$ (5)
By substituting the Rayleigh number prediction into (5), we calculate a flux Rayleigh number of $Ra_{F}\simeq 10^{28}$ in Earth’s core. Further substituting $k\simeq 100$ W/(mK) (Olson, 2015) and solving for $q_{conv}$ yields a prediction of $q_{conv}\simeq 3$ mW/m2 for the core. The superadiabatic heat flow is then calculated as $Q_{conv}=q_{conv}\cdot 4\pi r_{cmb}^{2}\simeq 0.5$ TW driving thermal convection in the core, where $r_{cmb}\simeq 3480$ km is the radius of the core mantle boundary. These predictive equations can be applied in other geophysical settings as well, such as subsurface oceans (e.g., Soderlund et al., 2024).

### v2

# Diffusivity-Free Turbulence in Liquid Metal Rotating Rayleigh-Bénard Convection Experiments

@@
Abstract
@@
Convection in planets and stars is predicted to occur in the “ultimate regime” of diffusivity-free, rapidly rotating turbulence, in which flows are characteristically unaffected by viscous and thermal diffusion. Boundary layer diffusion, however, has historically hindered experimental study of this regime. Here, we utilize the boundary-independent oscillatory thermal-inertial mode of rotating convection to realize the diffusivity-free scaling in liquid metal laboratory experiments. This oscillatory style of convection arises in rotating liquid metals (low Prandtl number fluids) and is driven by the temperature gradient in the fluid bulk, thus remaining independent of diffusive boundary dynamics. We triply verify the existence of the diffusivity-free regime via measurements of heat transfer efficiency $Nu$, dimensionless flow velocities $Re$, and internal temperature anomalies $\theta$, all of which are in quantitative agreement with planar asymptotically-reduced models. Achieving the theoretical diffusivity-free scalings in desktop-sized laboratory experiments provides the validation necessary to extrapolate and predict the convective flows in remote geophysical and astrophysical systems.
preprint: APS/123-QED
Figure 1: Rotating laboratory setup. (a) Image of the RoMag device at UCLA with the $\Gamma\simeq 1/2$ tank in place. Heat is supplied to the base (‘heating block’) by a heating pad and extracted from the lid (‘cooling block’) by a heat exchanger routed to a recirculating chiller. The cell is filled with liquid gallium and wrapped in layers of insulation (not pictured). Red and black visible wires connect to thermistors and Doppler transducer probes. (b) Schematic of measurement locations. White lines at the cell boundaries represent thermistors embedded within the heating and cooling blocks. Black dots indicate the location of five internal thermistors. Orange lines indicate the 1D path sensed by Doppler transducers. (c) Example time series data for a turbulent case in the oscillatory regime. The top panel shows thermistor data, where shades of pink (blue) indicate bottom (top) boundary temperatures and shades of gray indicate internal temperatures. The bottom panel shows velocimetry data, where pink (blue) regions indicate upward (downward) flow. The middle 50% of the chord is used to calculate the rms-velocity (demarcated with dotted lines). The case shown corresponds to parameters: $Ra\simeq 10^{9}$, $E\simeq 8\times 10^{-7}$, $Ro_{c}\simeq 0.16$, $Ro_{z}\simeq 0.01$, and $Re_{z}\simeq 2\times 10^{4}$.
@@
Diffusivity-free (DF) turbulent rotating convection is theorized to occur in the convection zones of stars and the fluid cores of planets, where it drives vortical flows, jets, and the dynamo action that sustains global-scale magnetic fields [1]. The DF regime is characterized by strong turbulence that is dominated by fluid inertia, and is unaffected by the fluid’s micro-scale viscosity ($\nu$) and thermal diffusivity ($\kappa$). Fluid physics researchers have long sought to validate this theoretical regime through measurements of heat transfer efficiency in numerical simulations and laboratory experiments [e.g., 2, 3]. It remains an essential question whether the theorized DF regime can be attained in a laboratory setting and extrapolated to natural settings.
@@
The canonical system for studying rotationally-constrained turbulent flows is the rotating Rayleigh-Bénard convection (RRBC) model [3]. In RRBC, a fluid layer confined between two parallel no-slip boundaries is heated from below and cooled from above while rotating about its central vertical axis. Scaling laws that relate global heat transport (Nusselt number, $Nu$) and bulk flow velocity (Reynolds number, $Re$) to the buoyancy forcing (Rayleigh number, $Ra$), global rotation period (Ekman number, $E$), and material properties (Prandtl number, $Pr$) enable the extrapolation of laboratory-numerical findings to more extreme geophysical and astrophysical scales [4, 5, 6, 7, 8, 9]. (See Table 1 for parameter definitions.)
@@
Rayleigh $Ra=\alpha g\Delta TH^{3}/(\nu\kappa)$ $\rm{\frac{thermal\ buoyancy}{diffusion}}$
@@
The scaling predictions for the DF regime are derived by enforcing the constraint that thermal and viscous diffusion have no effect on the total heat transported through the system nor the internal flow velocities, and assuming that the system’s angular rotation rate goes to infinity ($\Omega\rightarrow\infty$) [10, and see Supplementary Material]. The global heat transport prediction for rotating DF convection follows
@@
where $C_{J}$ is a constant prefactor. This value was determined by Julien et al. [10] to be $C_{J}\approx 1/25$ in doubly-periodic planar models of the asymptotically reduced RRBC equations across a range of low to moderate $Pr$ values. Scaling predictions for the internal quantities are determined from a torque balance between the inertial and buoyancy terms of the vorticity equation describing rotating flows, neglecting viscous diffusion [e.g., 11, and see Supplementary Material]. Assuming any scaling coefficients other than $C_{J}$ are order unity, the prediction for the internal flow velocity is
@@
where $\theta$ is the root-mean-squared temperature perturbation, and $Ro_{c}=(RaE^{2}Pr^{-1})^{1/2}$ is the convective Rossby number [9]. These scalings have yet to be unambiguously found in a closed cell RRBC laboratory setup, thus raising the question of their applicability to natural enclosed systems [cf. 12].
@@
The fundamental barrier to achieving diffusion-free convection in RRBC lies in the physical boundaries through which heat is supplied and extracted. Boundary layers throttle energy transfer through the system. In most RRBC cases, the primary mode of convective instability is that of steady convection, which is thermally controlled by the conditions at the boundaries [13, 14]. The critical value of the Rayleigh number for which steady convective motions first develop (denoted as $Ra_{S}$) is determined from a balance between friction and the Coriolis and pressure gradient forces, such that the steady mode is considered a viscous convective instability [15, 16]. Consequently, heat transfer in typical steady RRBC is dominated by boundary layer diffusion [e.g., 9, 17], which impedes the system from reaching the asymptotic DF scaling limit.
@@
Lateral boundaries in RRBC give rise to additional sidewall-attached convective instabilities. These “wall modes” manifest in the form of flow structures that travel azimuthally around the periphery of the tank, and are also a viscous instability [18]. Between contributions from steady bulk convection and wall mode convection, RRBC heat transfer is often strongly influenced by diffusive effects.
@@
In order to mitigate boundary contributions in convection studies, researchers have employed innovative techniques to bypass the boundaries altogether. Numerically, this has been achieved through the use of ghost layers or forced decoupling of boundary layers [19, 20]. Experimentally, radiative heat deposition directly into the fluid bulk [21, 22] has recently found the asymptotic heat transfer scaling exponent, but a very different value of the prefactor than that of Julien et al. [10] due to the differing means of thermal energy injection. Their results may apply to the rotating convective turbulence present in gas planets and stars; however, it is uncertain that they apply to terrestrial fluid environments where convection is still driven by fluxing heat through finite, solid boundaries. Thus, bounded RRBC systems have yet to yield unambiguous asymptotic scalings, raising the dual concerns $i$) that standard laboratory-scale devices may be unable to realistically simulate geophysical turbulence or $ii$) that DF scalings do not exist in natural convection enclosed by solid boundaries [23, 24, 25, 26, 27].
@@
To address these questions, we utilize the thermal-inertial oscillatory mode of convection that exists in low $Pr$ liquid metal RRBC [28, 29, 30, 31, 16]. Contrary to steady convection, which is subject to thermal boundary conditions, oscillatory convection is driven entirely by the internal temperature field. Zhang and Roberts [29] showed that oscillatory convection is described by a thermal-inertial wave, arising from a balance between inertia and Coriolis acceleration, with no dependence on the nature of the boundaries or diffusive effects at dominant order. This thermally-driven inertial instability arises exclusively in rotating low $Pr<0.68$ fluids, like liquid metals and plasmas. The critical $Ra$ for which oscillatory modes onset (denoted as $Ra_{O}$), occurs at lower $Ra$ than that for steady convection ($Ra_{S}$), such that thermal-inertial oscillations are considered the preferred mode of convection when the Prandtl number is sufficiently small [15]. Here, by isolating the oscillatory mode in rotating liquid gallium ($Pr\approx 0.027$) convection experiments, we show that the inertial asymptotic DF scalings (1) - (3) can be realized in a desktop-sized RRBC cell, and are in good quantitative agreement with Julien et al. [10]’s asymptotically-reduced low to moderate $Pr$ RRBC models, further validating both laboratory and multi-scale asymptotic methods.
@@
Figure 2: Rotating convection survey measurements. (a) Global heat transport efficiency, $Nu$, and (b) vertical non-dimensional bulk velocity, $Re_{z}$, as a function of thermal buoyancy forcing, $Ra$. Symbol color denotes the Ekman number, $E$, where red (lowest $E$) indicates the strongest Coriolis forcing and purple (highest $E$) indicates weaker Coriolis forcing. Symbol shape denotes the cell aspect ratio, $\Gamma=D/H$. Laboratory results from this study ($\Gamma\simeq 1/2$) are marked as circles, those from King and Aurnou [32] ($\Gamma\simeq 1$) are marked as triangles, and those from Vogt et al. [33] ($\Gamma\simeq 2$) are marked as squares. Hollow points indicate cases in which the steady mode of convection has onset ($Ra\geq Ra_{S}$), which is used here as a transition point between inertial and steady convection regimes. The solid black line in each panel shows the non-rotating empirically-determined scaling for the $\Gamma\simeq 1/2$ cell from this study for reference.
@@
UCLA’s RoMag device is used to conduct our experiments. The setup (Fig. 1a) features a right cylindrical convection cell with a height of $H=40.0$ cm and a diameter of $D=19.9$ cm, yielding an aspect ratio of $\Gamma=D/H\simeq 1/2$. The cell is constructed with a stainless steel sidewall positioned between two copper end blocks. Heat is supplied through a basal heating pad, while a heat exchanger plate, connected to a re-circulating chiller, removes heat at the top. After equilibration, both the top and bottom copper end blocks are maintained at fixed temperatures for the duration of the experiment. The convection cell is affixed to a rotating platform that spins the cell about its vertical axis. We carry out a systematic survey, varying the Rayleigh and Ekman numbers by approximately two orders of magnitude each ($10^{7}\leq Ra\leq 10^{9}$ and $3\times 10^{-7}\leq E\leq 10^{-5}$).
@@
For each experiment, we acquire measurements of temperature, heat flux, and velocity. Thermistors embedded within the top and bottom thermal blocks measure the vertical temperature gradient, $\Delta T$ (see white lines in Fig. 1b and pink/blue time series lines in Fig. 1c). Five internal thermistors are positioned within the bulk fluid to measure the internal temperature fluctuations (see black lines in Fig. 1b and black/gray time series lines in Fig. 1c). The standard deviation of each internal thermistor time series is calculated to determine the local perturbation at each location, which are then RMS-averaged to yield $\theta$. Vertical flow velocity, $u_{z}$, is measured using an ultrasonic Doppler velocimeter (UDV). A Doppler transducer probe is positioned in contact with the interior fluid, oriented along the vertical axis of the cylinder. Natural impurities in the gallium act as particles that the UDV can acoustically sense [34]. Velocities are recorded at approximately 2 mm increments across the entire vertical length of the cell at a sampling rate of approximately 10 Hz. The characteristic $u_{z}$ velocity is then calculated as a time- and space-RMS, using the middle 50% of the cell height (see orange lines and light orange area shaded in Fig. 1b and the Hovmöller diagram in 1c).
@@
We compare our $\Gamma\simeq 1/2$ convection cell results to those of two other aspect ratios from prior studies performed using the same experimental device. King and Aurnou [32] conducted rotating convection experiments in liquid metal using a $\Gamma\simeq 1$ cell, measuring global heat transport efficiency, $Nu$, with their data reported in King and Aurnou [35]. Vogt et al. [33] conducted experiments using a $\Gamma\simeq 2$ cell, measuring the Nusselt number, $Nu$, vertical Reynolds number, $Re_{z}$, and internal temperature fluctuations, $\theta$. We additionally compare our findings to the asymptotically-reduced planar models of Julien et al. [10] by incorporating their best-fit prefactor $C_{J}=1/25$ in the heat transfer prediction (1). We elect to carry this prefactor through the asymptotic predictions for the Reynolds number (2) and thermal anomalies (3), as shown in the Supplementary Material.
@@
Figure 3: Laboratory scaling results. (a) Convective heat flux relative to conductive heat flux, $Nu-1$, (b) normalized internal temperature perturbation, $\theta/\Delta T$, and (c) measured vertical Reynolds number, $Re_{z}$, each shown as a function of the asymptotically reduced scaling result from Julien et al. [10]. Solid black lines indicate the asymptotically reduced predictions. Dotted lines show the best-fit power law results to each $\Gamma$ data set, where the hollow points (indicating $Ra/Ra_{S}\geq 1$) are excluded from the fit. Inset plots display the best-fit prefactors as hollow symbols and the best fit scaling exponents as solid symbols, with solid horizontal lines marking the asymptotic predictions. The best-fit scaling relations are each listed in Table 2. Across all three measured quantities, the scaling approaches the asymptotic result as the aspect ratio, $\Gamma$, increases. Other markers are the same as those in Fig. 2.
@@
Fig. 2 presents the measured heat transfer ($Nu$) and vertical velocities ($Re_{z}$) across all three aspect ratios considered, plotted as a function of the buoyancy forcing ($Ra$). Rotating data are indicated by colored markers, while a solid black line represents the empirically determined non-rotating scaling relation. Solid colored points indicate data within the oscillatory convection regime. Hollow points represent experiments where steady convection has additionally onset, following $Ra\geq Ra_{S}=(8.70-9.63E^{1/6})E^{-4/3}$ [36, 37]. In the parameter ranges investigated here, these steady cases are not expected to follow the inertial DF scaling predicted for the oscillatory mode and are therefore not included in the scaling analysis. (Under realistic geophysical conditions, [10] predict that steady rotating convective modes will also exist in the DF regime.) In addition, a weakly nonlinear regime is present near onset ($1\leq Ra/Ra_{O}\lesssim 1.5$) [38, 39]. Thus, only measurements where $Ra/Ra_{O}>1.5$ are considered here. Notably, all cases exhibit high Reynolds numbers with $300\lesssim Re\lesssim 3000$. Even near convective onset ($Nu\leq 2$), the Reynolds numbers reach as high as $Re\simeq 10^{3}$.